# Cascade adaptive optics with a second stage based on a Zernike wavefront sensor for exoplanet observations

### v1

@@ Abstract
Context. For the past decade, the high-contrast observation of disks and gas giant planets around nearby stars has been made possible on ground-based instruments using extreme adaptive optics (XAO). While these facilities produce images with a Strehl ratio larger than 90% in H-band in median observing conditions and high-flux regime, the correction leaves adaptive optics (AO) residuals which prevent the study of fainter or less massive exoplanets.
Aims. Cascade AO systems with a fast second stage based on a Pyramid wavefront sensor (PWFS) have recently emerged as an appealing solution to reduce the atmospheric wavefront errors. Since these phase aberrations are expected to be small, they can also be accurately measured by a Zernike wavefront sensor (ZWFS), a well-known concept for its high sensitivity and moderate linear capture range. We propose an alternative second stage relying on the ZWFS to correct for the AO residuals.
Methods. The cascade AO with a ZWFS-based control loop is implemented on the ESO’s GPU-based High-order adaptive OpticS Testbench (GHOST) to validate the scheme in monochromatic light. We emulate the XAO first stage in different observing conditions (wind speed, seeing) and determine corresponding operation parameters (e.g., number of controlled modes, integrator gain, loop calibration) that lead to stable loop operation and good correction performance. Our strategy is assessed in terms of corrected wavefront errors and contrast gain in the images with a Lyot coronagraph to probe its efficiency.
Results. In median wind speed and seeing, our second-stage AO with a ZWFS and a basic integrator reduce the atmospheric residuals by a factor of 6 and increase the wavefront error stability with a gain of 2 from open to closed loop. In the presence of non common path aberrations, we also reach a contrast gain by a factor of 2 in the coronagraphic images at short separations from the source, proving the ability of our scheme to work in cascade with an XAO loop and help imaging fainter or lighter close-in companions. In more challenging conditions, contrast improvements are also achieved by adjusting the control loop features.
Conclusions. Our study validates the ZWFS-based second-stage AO loop as an effective solution to address small residuals left from a single-stage XAO system for the coronagraphic observations of circumstellar environments. Our first in-lab demonstration paves the way for more advanced versions of our approach with different temporal control laws, non-linear reconstructors, and spectral widths to operate with high-contrast facilities on the current 8-10 m class telescopes and Extremely Large Telescopes and observe planets down to Earth analogs around M dwarfs.
@@
Unveiling the mechanisms of formation, evolution, and habitability of exoplanetary systems is one of the most exciting topics in modern astrophysics. The discovery of over 5,500 exoplanets to date111e.g. see https://exoplanet.eu/home/ or https://exoplanetarchive.ipac.caltech.edu/ with different detection methods has enriched our understanding of the demography and diversity of planetary systems, the morphology of circumstellar disks, and the internal structure and atmosphere of planets. High-contrast observation is a direct detection method that allows us to study the outer part of exoplanetary systems beyond the 5-10 AU range, explore interactions between planets and disks, and characterize the atmospheric properties of exoplanets (e.g. Currie et al. 2023; Chauvin 2024).
@@
However, extracting the photons from exoplanetary companions is extremely challenging, given the contrast, i.e. the flux ratio between a star and a planet, at short angular separation in the visible and near infrared (Oppenheimer & Hinkley 2009; Traub & Oppenheimer 2010). Nowadays, the foremost high-contrast facilities on the ground (e.g. Gemini/GPI, Subaru/SCExAO, Keck/KPIC, VLT/SPHERE, Magellan/MagAO-X, Macintosh et al. 2014; Jovanovic et al. 2015; Mawet et al. 2016; Beuzit et al. 2019; Males et al. 2022) or in space (e.g. HST, JWST, Debes et al. 2019; Carter et al. 2021; Hinkley et al. 2022) observe gas giant planets which are 104 to 106 times fainter than their stellar hosts at angular separations down to 200 mas in emitted light. In the forthcoming years, some of these exoplanet imagers will benefit from upgrades to image and spectrally analyze Jovian and Neptune-size planets which are up to 108 times fainter than their host star at angular separations down to 100 mas in thermal and reflected light (Jovanovic et al. 2019; Ahn et al. 2021; Boccaletti et al. 2022; Chilcote et al. 2022; Males et al. 2022). In the next decades, the Extremely Large Telescopes on the ground (ELTs, Kasper et al. 2021; Fitzgerald et al. 2022; Kautz et al. 2023) or observatories in space (e.g. Roman Space Telescope, LUVOIR/HabEx mission concepts, Bailey et al. 2023; The LUVOIR Team 2019; Gaudi et al. 2020) will benefit from high-contrast capabilities to study small gaseous and rocky planets that are 108 to 1010 fainter than their parent stars at an angular separation of less than 20 mas in reflected light.
@@
High-contrast observations rely on different features such as extreme adaptive optics (XAO), wavefront sensing and control, stellar coronagraphy, observing strategies, and image post-processing methods. Instrumental and numerical innovations are required in all these aspects to gain up to four orders of magnitude in contrast for the observation of these faint, close-in planetary-mass companions (Galicher & Mazoyer 2024). The associated technological leaps will lead to a refined understanding of the formation of Jovian and terrestrial planets, the survey of Earth analogs around Sun-like stars, and the tentative detection of biosignatures from nearby habitable zone exoplanets by the 2040s.
@@
On the ground, the most advanced high-contrast facilities are built upon an XAO system to compensate for the effects of the atmospheric turbulence on the image of an observed star (Guyon 2018). Usually based on a Shack-Hartmann wavefront sensor (SHWFS) to measure disturbances, these systems run at a typical speed of 1 kHz to provide images with a Strehl ratio larger than 90% in H-band in median conditions and in high-flux regime (e.g. Perrin et al. 2003; Fusco et al. 2006; Beuzit et al. 2019). Following the Maréchal approximation, the corresponding wavefront errors are of the order of 85 nm Root Mean Square (RMS) or 0.05 at =1.65 m where denotes the wavelength of observation.
@@
While these XAO systems perform an excellent correction of the atmospheric wavefront perturbations (e.g. Fusco et al. 2016), they leave some adaptive optics (AO) residuals which prevent the observation of the substellar mass companions. Upgrading these key modules will pave the way for novel exoplanetary science capabilities, such as the access to young gas giant planets down to the snow line or the observation of a large number of red stars (Boccaletti et al. 2020). Observing fainter or closer planets to their star involves achieving image contrast larger than 106 down to 100 mas and gaining sensitivity for red stars.
@@
To reduce AO residuals, the XAO systems require a more sensitive wavefront sensor (WFS) than SHWFS and a larger temporal bandwidth than 1 kHz to achieve more accurate wavefront error measurements and faster corrections than the evolution of the atmospheric turbulence. Upgrading the existing XAO systems is an appealing solution but a risky option from the operational standpoint since the availability of the current functionalities of these exoplanet imagers is often requested by the community to preserve the existing high-contrast capabilities. To avoid modifying the initial XAO, a promising approach consists in using a cascade AO system (e.g. Cerpa-Urra et al. 2022). This concept consists in assisting the original XAO system with a subsequent AO stage including its own deformable mirror (DM), WFS and fast real-time computer (RTC) to decrease the atmospheric residuals further.
@@
The regime of expected atmospheric residuals also appears favorable to a second-stage AO loop with a Zernike wavefront sensor (ZWFS), a concept based on phase-contrast principle to convert wavefront errors into intensity variations (Zernike 1934). Such a sensor is well-known for its moderate capture range and its high sensitivity (e.g. Bloemhof & Wallace 2003; Dohlen 2004; Guyon 2005; Wallace et al. 2011; Jensen-Clem et al. 2012; N’Diaye et al. 2013; Chambouleyron et al. 2021). This concept proves a promising solution for a wide range of astronomical applications, such as the measurement of non-common path aberrations (NCPA) in high-contrast instruments (e.g. ZELDA on VLT/SPHERE, N’Diaye et al. 2016; Vigan et al. 2019, 2022), the estimate of cophasing errors in segmented aperture telescopes (Dohlen et al. 2006; Surdej et al. 2010; Vigan et al. 2011; Janin-Potiron et al. 2017; Pourcelot et al. 2021; van Kooten et al. 2022; Salama et al. 2024), the measurement of fine low-order aberrations in space observatories with coronagraphic capabilities (e.g. Shi et al. 2016; Pourcelot et al. 2022, 2023), or the picometric precision metrology for future space coronagraphs (e.g., Ruane et al. 2020; Steeves et al. 2020).
In this paper, we investigate a ZWFS-based approach to run fast and accurate wavefront corrections on the image of an observed star after an XAO loop. Our ZWFS-based AO loop is implemented on the GPU-based High-order adaptive OpticS Testbench (GHOST) at the ESO Headquarters in Garching (Germany) to demonstrate its efficiency in simulated median observing conditions. We consider different control loop aspects such as the number of corrected modes, the temporal loop gain, and the calibration to assess their impact on the wavefront correction. The ZWFS control loop is then explored through different observing conditions to derive the best functioning points. Finally, we perform some comparisons with the PWFS-based approach to derive the complementary working regimes for our ZWFS-based control loop. Our tests were run in Fall 2023. The whole study is carried out in the presence of a monochromatic light source to achieve a first experimental validation of our scheme.
@@ 2.1 GHOST testbed setup
GHOST is equipped with two different light sources: a light emitting diode (LED) and a super luminescent diode (SLED) emitting at 740nm and 770nm respectively. Our work here relies on the use of the SLED source. Based on a single-mode fiber, a point-like source emits a light beam which is collimated by the achromatic lens L1. The light then goes through a polarizing beam splitter cube, a standard beam splitter BS1 and a 10 mm-diameter diaphragm, leading to a polarized light beam that hits the Meadowlark liquid crystal on silicon (LCoS) spatial light modulator (SLM) in reflection to inject simulated XAO residual turbulence phase screens.
The reflected light from the SLM travels back to BS1 and is then resized with a beam expander to 6.7mm before hitting the second beam splitter BS2. The transmitted light goes towards the 492-1.5 DM from Boston Micromachine Corporation (BMC) with an actuator pitch and stroke of 300 m and 1.5 m. The DM is used as a corrector for the second-stage AO to reduce the XAO residuals from the SLM. After reflection, the beam travels back to BS2 and then faces the beam splitter BS3 which equally splits the light between the wavefront sensing and the science arms.
@@
The transmitted light after BS3 goes to the wavefront sensing arm with a lens that forms the source image on the sensor with a f/50 beam ratio. In this focal plane, a field stop can also be inserted to reduce aliasing effects in the WFS measurements (e.g. Poyneer & Macintosh 2004). Field stops with diameters of 35, 49 and 63 are available to adjust as the function of the atmospheric conditions. A downstream lens forms the image of the relayed pupil onto a 10GigE Sony IMX CMOS camera to capture the sensor signal for the wavefront error measurements. The pupil is imaged with 36 pixels across its diameter. The standard GHOST setup uses the PWFS to estimate the residual atmospheric aberrations. The wavefront sensing arm also includes an additional PI SL-325 modulation mirror to enable beam modulation on the PWFS. This device is not represented in the scheme. In this work, we replace the Pyramid with a Zernike phase-shifting mask.
@@ 2.2.1 Principle
We briefly recall the principle of the ZWFS, see its optical layout in Fig. 2. The light beam from a point-like source propagates through the sensor, first going through the telescope pupil in plane A in which phase aberrations are present. The source image is then formed in the following focal plane B where a mask with a phase shift of and a diameter of about a resolution element is located. The light going through and surrounding the mask dot generates interferences in the relayed pupil plane C, producing an intensity pattern that is directly related to . In recent years, the formalism of the ZWFS has been largely described for applications in astronomy, and further details can be found in the literature (e.g. N’Diaye et al. 2013; Ruane et al. 2020; Chambouleyron et al. 2021). Following N’Diaye et al. (2013), the relayed pupil intensity writes as a function of the phase for a given pixel in the pupil with
where and represent the amplitude pupil function and the amplitude diffracted by the focal plane phase mask. In the absence of aberrations (), the sensor response is non-null and its intensity writes as
In the small aberration regime ( 1 rad), we refer to the first-order Taylor expansion of the terms with to derive a linear relation from Eq. (1) between the intensity variations in the relayed pupil and the wavefront errors in the entrance pupil. Such an approach favors a simple and fast retrieval of small aberrations. Assuming that , , and are different from zero, the phase is then derived from with
@@
These expressions are valid for the phase aberrations which correspond to a range for the optical path differences. In this linear regime, the phase aberrations can alternatively be reconstructed with the ZWFS by relying on a command matrix built from an interaction matrix, following the standard approach in adaptive optics. In this paper, our phase retrieval method with ZWFS will be based on the use of interaction and command matrices.
@@ 2.2.2 GHOST mask prototype
The metrology of the prototype was carried out at Laboratoire d’Astrophysique de Marseille with a Wyko NT9100 interference microscope in vertical scanning mode, allowing us to confirm that the measured mask diameter and depth are within 1% of the specifications that were given to the manufacturer.
For our tests, the mask is mounted on xy-translation stages to adjust its position with respect to the source image.

### v2

@@ Abstract
Context. Over the past decade, the high-contrast observation of disks and gas giant planets around nearby stars has been made possible with ground-based instruments using extreme adaptive optics (XAO). These facilities produce images with a Strehl ratio higher than 90% in the H band, in median observing conditions and high-flux regime. However, the correction leaves behind adaptive optics (AO) residuals, which impede studies of fainter or less massive exoplanets.
Aims. Cascade AO systems with a fast second stage based on a Pyramid wavefront sensor (PWFS) have recently emerged as an appealing solution to reduce the atmospheric wavefront errors. Since these phase aberrations are expected to be small, they can also be accurately measured by a Zernike wavefront sensor (ZWFS), a well-known concept for its high sensitivity and moderate linear capture range. We propose an alternative second stage that relies on the ZWFS to correct for the AO residuals.
Methods. We implemented the cascade AO with a ZWFS-based control loop on the ESO’s GPU-based High-order adaptive OpticS Testbench (GHOST) to validate the scheme in monochromatic light. We emulated the XAO first stage in different observing conditions (wind speed, seeing) and determined the corresponding operation parameters (e.g., number of controlled modes, integrator gain, loop calibration) that lead to stable loop operation and good correction performance. Our strategy was assessed in terms of corrected wavefront errors and contrast gain in the images with a Lyot coronagraph to probe its efficiency.
Results. In median wind speed and seeing, our second-stage AO with a ZWFS and a basic integrator was able to reduce the atmospheric residuals by a factor of 6 and increase the wavefront error stability with a gain of 2 between open and closed loop. In the presence of non-common path aberrations, we also achieved a contrast gain of a factor of 2 in the coronagraphic images at short separations from the source, proving the ability of our scheme to work in cascade with an XAO loop. In addition, it may prove useful for imaging fainter or lighter close-in companions. In more challenging conditions, contrast improvements are also achieved by adjusting the control loop features.
Conclusions. Our study validates the ZWFS-based second-stage AO loop as an effective solution to address small residuals left over from a single-stage XAO system for the coronagraphic observations of circumstellar environments. Our first in-lab demonstration paves the way for more advanced versions of our approach with different temporal control laws, non-linear reconstructors, and spectral widths. This would allow our approach to operate in high-contrast facilities on the current 8-10 m class telescopes and Extremely Large Telescopes to observe exoplanets, all the way down to Earth analogs around M dwarfs.
@@
Unveiling the mechanisms of formation, evolution, and habitability of exoplanetary systems is one of the most exciting topics in modern astrophysics. The discovery of over 5500 exoplanets to date111e.g., see https://exoplanet.eu/home/ or https://exoplanetarchive.ipac.caltech.edu/ with different detection methods has enriched our understanding of the demography and diversity of planetary systems, the morphology of circumstellar disks, and the internal structure and atmosphere of planets. High-contrast observation is a direct detection method that allows us to study the outer part of exoplanetary systems beyond the 5-10 AU range, explore interactions between planets and disks, and characterize the atmospheric properties of exoplanets (e.g., Currie et al. 2023; Chauvin 2024).
@@
However, extracting the photons from exoplanetary companions is extremely challenging, given the contrast, namely, the flux ratio between a star and a planet, at short angular separation in the visible and near infrared (Oppenheimer & Hinkley 2009; Traub & Oppenheimer 2010). Nowadays, the foremost high-contrast facilities on the ground (e.g., Gemini/GPI, Subaru/SCExAO, Keck/KPIC, VLT/SPHERE, Magellan/MagAO-X, Macintosh et al. 2014; Jovanovic et al. 2015; Mawet et al. 2016; Beuzit et al. 2019; Males et al. 2022) or in space (e.g., HST, JWST, Debes et al. 2019; Carter et al. 2021; Hinkley et al. 2022) observe gas giant planets which are 104 to 106 times fainter than their stellar hosts at angular separations down to 200 mas in emitted light. In the coming years, some of these exoplanet imagers will benefit from upgrades to image and spectrally analyze Jovian and Neptune-size planets, which are up to 108 times fainter than their host star at angular separations down to 100 mas in thermal and reflected light (Jovanovic et al. 2019; Ahn et al. 2021; Boccaletti et al. 2022; Chilcote et al. 2022; Males et al. 2022). Over the next few decades, the Extremely Large Telescopes on the ground (ELTs, Kasper et al. 2021; Fitzgerald et al. 2022; Kautz et al. 2023) or observatories in space (e.g., Roman Space Telescope, LUVOIR/HabEx mission concepts, Bailey et al. 2023; The LUVOIR Team 2019; Gaudi et al. 2020) will benefit from high-contrast capabilities to study small gaseous and rocky planets that are 108 to 1010 fainter than their parent stars at an angular separation of less than 20 mas in reflected light.
@@
High-contrast observations rely on different features such as extreme adaptive optics (XAO), wavefront sensing and control, stellar coronagraphy, observing strategies, and image post-processing methods. Instrumental and numerical innovations are required in all these aspects to gain up to four orders of magnitude in terms of contrast for the observation of these faint, close-in planetary-mass companions (Galicher & Mazoyer 2024). The associated technological leaps will lead to a refined understanding of the formation of Jovian and terrestrial planets, the survey of Earth analogs around Sun-like stars, and the tentative detection of biosignatures from nearby habitable zone exoplanets by the 2040s.
@@
On the ground, the most advanced high-contrast facilities are built upon an XAO system to compensate for the effects of the atmospheric turbulence on the image of an observed star (Guyon 2018). Usually based on a Shack-Hartmann wavefront sensor (SHWFS) to measure disturbances, these systems run at a typical speed of 1 kHz to provide images with a Strehl ratio higher than 90% in H-band in median conditions and in high-flux regime (e.g., Perrin et al. 2003; Fusco et al. 2006; Beuzit et al. 2019). Following the Maréchal approximation, the corresponding wavefront errors are of the order of 85 nm Root Mean Square (RMS) or 0.05 at =1.65 m, where denotes the wavelength of observation.
@@
While these XAO systems perform an excellent correction of the atmospheric wavefront perturbations (e.g., Fusco et al. 2016), they leave behind some adaptive optics (AO) residuals, which impede the observation of the substellar mass companions. Upgrading these key modules will pave the way for novel exoplanetary science capabilities, such as access to young gas giant planets down to the snow line or the observation of a large number of red stars (Boccaletti et al. 2020). Observing fainter or closer planets to their star involves achieving image contrast larger than 106 down to 100 mas and gaining sensitivity for red stars.
@@
To reduce AO residuals, the XAO systems require a more sensitive wavefront sensor (WFS) than SHWFS and a higher temporal bandwidth than 1 kHz to achieve more accurate wavefront error measurements and faster corrections than the evolution of the atmospheric turbulence. Upgrading the existing XAO systems is an appealing solution but a risky option from the operational standpoint since the availability of the current functionalities of these exoplanet imagers is often requested by the community to preserve the existing high-contrast capabilities. To avoid modifying the initial XAO, a promising approach consists of using a cascade AO system (e.g., Cerpa-Urra et al. 2022). This concept consists of assisting the original XAO system with a subsequent AO stage by including its own deformable mirror (DM), WFS, and fast real-time computer (RTC) to decrease the atmospheric residuals further.
@@
The regime of expected atmospheric residuals also appears favorable to a second-stage AO loop with a Zernike wavefront sensor (ZWFS), a concept based on phase-contrast principle to convert wavefront errors into intensity variations (Zernike 1934). Such a sensor is well-known for its moderate capture range and its high sensitivity (e.g., Bloemhof & Wallace 2003; Dohlen 2004; Guyon 2005; Wallace et al. 2011; Jensen-Clem et al. 2012; N’Diaye et al. 2013; Chambouleyron et al. 2021). This concept has proven to be a promising solution for a wide range of astronomical applications, such as the measurement of non-common path aberrations (NCPA) in high-contrast instruments (e.g., ZELDA on VLT/SPHERE, N’Diaye et al. 2016; Vigan et al. 2019, 2022), the estimate of cophasing errors in segmented aperture telescopes (Dohlen et al. 2006; Surdej et al. 2010; Vigan et al. 2011; Janin-Potiron et al. 2017; Pourcelot et al. 2021; van Kooten et al. 2022; Salama et al. 2024), measurement of fine low-order aberrations in space observatories with coronagraphic capabilities (e.g., Shi et al. 2016; Pourcelot et al. 2022, 2023), or the picometric precision metrology for future space coronagraphs (e.g., Ruane et al. 2020; Steeves et al. 2020).
In this paper, we investigate a ZWFS-based approach to run fast and accurate wavefront corrections on the image of an observed star after an XAO loop. Our ZWFS-based AO loop is implemented on the GPU-based High-order adaptive OpticS Testbench (GHOST) at the ESO Headquarters in Garching (Germany) to demonstrate its efficiency in simulated median observing conditions. We consider different control loop aspects such as the number of corrected modes, the temporal loop gain, and the calibration to assess their impact on the wavefront correction. The ZWFS control loop is then explored through different observing conditions to derive the best functioning points. Finally, we perform some comparisons with the PWFS-based approach to derive the complementary working regimes for our ZWFS-based control loop. Our tests were run in autumn 2023. The whole study is carried out in the presence of a monochromatic light source to achieve a first experimental validation of our scheme.
@@ 2.1 GHOST testbed setup
GHOST is equipped with two different light sources: a light emitting diode (LED) and a super luminescent diode (SLED) emitting at 740nm and 770nm, respectively. Our work here relies on the use of the SLED source. Based on a single-mode fiber, a point-like source emits a light beam which is collimated by the achromatic lens L1. The light then goes through a polarizing beam splitter cube, a standard beam splitter BS1 and a 10 mm-diameter diaphragm, leading to a polarized light beam that hits the Meadowlark liquid crystal on silicon (LCoS) spatial light modulator (SLM) in reflection to inject simulated XAO residual turbulence phase screens.
The reflected light from the SLM travels back to BS1 and is then resized with a beam expander to 6.7mm before hitting the second beam splitter BS2. The transmitted light goes towards the 492-1.5 DM from Boston Micromachine Corporation (BMC) with an actuator pitch and stroke of 300 m and 1.5 m. The DM is used as a corrector for the second-stage AO to reduce the XAO residuals from the SLM. After reflection, the beam travels back to BS2 and then faces the beam splitter BS3, which splits the light between the wavefront sensing and the science arms equally.
@@
The transmitted light after BS3 goes to the wavefront sensing arm with a lens that forms the source image on the sensor with a f/50 beam ratio. In this focal plane, a field stop can also be inserted to reduce aliasing effects in the WFS measurements (e.g., Poyneer & Macintosh 2004). Field stops with diameters of 35, 49 and 63 are available to adjust as the function of the atmospheric conditions. A downstream lens forms the image of the relayed pupil onto a 10GigE Sony IMX CMOS camera to capture the sensor signal for the wavefront error measurements. The pupil is imaged with 36 pixels across its diameter. The standard GHOST setup uses the PWFS to estimate the residual atmospheric aberrations. The wavefront sensing arm also includes an additional PI SL-325 modulation mirror to enable beam modulation on the PWFS. This device is not represented in the scheme. In this work, we replaced the Pyramid with a Zernike phase-shifting mask.
@@ 2.2.1 Principle
We briefly recall the principle of the ZWFS, see its optical layout in Fig. 2. The light beam from a point-like source propagates through the sensor, first going through the telescope pupil in plane A in which phase aberrations are present. The source image is then formed in the following focal plane B, where a mask with a phase shift of and a diameter, , of about one resolution element is located. The light going through and surrounding the mask dot generates interferences in the relayed pupil plane C, producing an intensity pattern that is directly related to . In recent years, the formalism of the ZWFS has been largely described for applications in astronomy and further details can be found in the literature (e.g., N’Diaye et al. 2013; Ruane et al. 2020; Chambouleyron et al. 2021). Following N’Diaye et al. (2013), the relayed pupil intensity is expressed as a function of the phase for a given pixel in the pupil with
where and represent the amplitude pupil function and the amplitude diffracted by the focal plane phase mask. In the absence of aberrations (), the sensor response is non-null and its intensity, , is expressed as
In the small aberration regime ( 1 rad), we refer to the first-order Taylor expansion of the terms with to derive a linear relation from Eq. (1) between the intensity variations in the relayed pupil and the wavefront errors in the entrance pupil. Such an approach favors a simple and fast retrieval of small aberrations. Assuming that , , and are not equal to zero, the phase is then derived from with
@@
These expressions are valid for the phase aberrations which correspond to a range for the optical path differences. In this linear regime, the phase aberrations can alternatively be reconstructed with the ZWFS by relying on a command matrix built from an interaction matrix, following the standard approach in AO. In our approach, the phase retrieval method with ZWFS is based on the use of interaction and command matrices.
@@ 2.2.2 GHOST mask prototype
The metrology of the prototype was carried out at Laboratoire d’Astrophysique de Marseille with a Wyko NT9100 interference microscope in vertical scanning mode, allowing us to confirm that the measured mask diameter and depth are within 1% of the specifications that were given to the manufacturer. For our tests, the mask is mounted on xy-translation stages to adjust its position with respect to the source image.